a tugboat is 36km due north of lighthouse C. Lighthouse B is directly east of lighthouse C. The lighthouses are 53km apart. Find the bearing og lighthouse B from the tugboat and the distance of lighthouse B from the tugboat.


1
Expert's answer
2021-04-27T11:44:54-0400

The figure of this problem is



From this figure,


"tan\alpha = \dfrac{53}{36}"


"\alpha = tan^{-1}(\dfrac{53}{36})"


"\alpha = 55.81 ^ \circ"

Now, we can use Pythagorean Theorem


"d = \sqrt{53^2+36^2}"

"d = 64.07 km"